OpenStudy (anonymous):

How do you find the integral of (sqrt(cot x)(csc^2 x) dx ?

OpenStudy (anonymous):

\[\int\sqrt{\cot x}~\csc^2x~dx\] Let \(u=\cot x\), then \(du=-\csc^2x~dx\), or \(-du=\csc^2x~dx\). So you have \[-\int\sqrt u~du\]

OpenStudy (anonymous):

I'm still confused about how o find it could you explain a little more?

OpenStudy (johnweldon1993):

\[\large \int \sqrt{\color \green{cot(x)}} \color \red{csc^2(x)dx}\] you'll see why I highlighted ina bit We do a u-substitution letting u = cot(x) We know the derivative of cot(x) = -csc^2(x) so \[\large u = \color \green{cot(x)}\]\[\large du = \color \red{-csc^2(x)dx}\] So what we are left with is (look at the relation between the highlights) \[\large \int- \sqrt{u}du\] We dont need that '-' sign right now...so lets just factor it out \[\large - \int \sqrt{u} du\] Now what is the integral of √u? We know that \(\large \sqrt{x} = x^{1/2}\) so \[\large - \int u^{1/2}du\] Now just using basic integration steps we know that \[\large \int x^a = \frac{x^{a + 1}}{a + 1}\] so \[\large - \int \frac{u^{1/2 + 1}}{1/2 + 1} = -\frac{u^{3/2}}{3/2} = -\frac{2u^{3/2}}{3}\] And lastly, we must plug back in 'u' in terms of 'x'...we originally made u = cot(x) so \[\large -\frac{2cot(x)^{3/2}}{3} + c\]

OpenStudy (anonymous):

Thank you so much!!! I understand it now(:

OpenStudy (johnweldon1993):

Anytime :)